In 2017 a tax payer received $1000 of tips that should of been reported to the employer, but did ...

in 2017 a tax payer received $1000 of tips that should of been reported to the employer, but did not. other income for the year was $20000. how much additional social security tax will the taxpayer pay when she completes form 4137.
Answer #1

Social Security Tax rate in United States is 6.2%

Since $1000 was not Included in Taxable Income so Social Security Tax will be 6.2% of $1000 = (1000 x 6.2)/100 = $62.

Social Security tax on other Income $20000 is 6.2/100*20000 = $1240

So Total Additional tax to be paid is 1240+62 = 1302$